# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ Appendix D $-\ell$ Achievability Analysis for Bernoulli Designs
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime}}\geq(1-2\zeta-o(1))\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}\geq\frac{d_{0}\cdot\rho\nu\ell ne^{-\nu}}{k}\Big{)}$ (121)
$\quad\quad+\sum_{\begin{subarray}{c}0\leq d<d_{0}\\ \frac{d\rho\nu e^{-\nu}\ell n}{k}\in\mathbb{Z}\end{subarray}}\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}=\frac{d\cdot\rho\nu\ell ne^{-\nu}}{k}\Big{)}\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},2}\leq\big{[}\rho d-(1-2\zeta-o(1))C\big{]}\frac{n\nu e^{-\nu}\ell}{k}\Big{)}.$ (124)
To bound the terms in (121) and (124), we consider the following two cases.
@@
Bounding the term in (121): Consider conditioning on $(\mathbf{X}_{s},\mathbf{Y})$ and utilizing the randomness of $\mathbf{X}_{\mathcal{J}^{\prime}}$, and recall from (119) and (105) that we have $\widetilde{G}_{\mathcal{J}^{\prime},1}\sim\text{Bin}(\rho ne^{-\nu}(1+o(1)),\frac{\nu\ell(1+o(1))}{k})$. Since $d_{0}>1$, we can apply Chernoff bound ((23) in Lemma 1) to bound the term in (121) as
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}\geq\frac{d_{0}\cdot\rho\nu\ell ne^{-\nu}}{k}\Big{)}$ $\leq\exp\left(-\rho ne^{-\nu}(1+o(1))D\Big{(}\frac{d_{0}\ell\nu}{k}\big{\|}\frac{\ell\nu}{k}\Big{)}\right).$ (125)
Bounding the term in (124): First observe that $\mathbb{P}(\widetilde{G}_{\mathcal{J}^{\prime},2}\leq[\rho d-(1-2\zeta-o(1))C]\frac{n\nu e^{-\nu}\ell}{k})=0$ if $d<\frac{C(1-2\zeta)}{\rho}-o(1)$, and so we can further restrict the summation over $d$ in (124) to the range
$\mathcal{D}:=\Big{\{}\max\Big{\{}0,\frac{C(1-2\zeta)}{\rho}\Big{\}}\leq d<d_{0}:\frac{d\rho\nu e^{-\nu}\ell n}{k}\in\mathbb{Z}\Big{\}}$ (126)
up to a $o(1)$ term that has no impact on our subsequent analysis. Since $d_{0}=O(1)$, the term in (124) involves no more than $O(\frac{\ell n}{k})$ summands, so using the Chernoff bound (see (23) in Lemma 1) we can bound it as
$O\Big{(}\frac{\ell n}{k}\Big{)}\max_{d\in\mathcal{D}}\Big{\{}\exp\Big{(}-\rho ne^{-\nu}(1+o(1))\cdot D\Big{(}\frac{d\nu\ell(1+o(1))}{k}\big{\|}\frac{\nu\ell(1+o(1))}{k}\Big{)}\Big{)}$ (127)
$\quad\quad\quad\cdot\exp\Big{(}-(1-\rho)ne^{-\nu}(1+o(1))\cdot D\Big{(}\frac{[\rho d-(1-2\zeta)C]\nu\ell(1+o(1))}{(1-\rho)k}\big{\|}\frac{\nu\ell(1+o(1))}{k}\Big{)}\Big{)}\Big{\}}$ (128)
$=\exp\Big{(}-ne^{-\nu}(1+o(1))\min_{d\in\mathcal{D}}\Big{[}\rho\cdot D\Big{(}\frac{d\nu\ell}{k}\big{\|}\frac{\nu\ell}{k}\Big{)}+(1-\rho)D\Big{(}\frac{[\rho d-(1-2\zeta)C]\nu\ell}{(1-\rho)k}\big{\|}\frac{\nu\ell}{k}\Big{)}\Big{]}\Big{)}$ (129)
$\leq\exp\Big{(}-\frac{\ell n\nu e^{-\nu}}{k}\big{[}\min_{d\geq\max\{0,\frac{A}{\rho}\}}g(C,\zeta,d,\rho)+o(1)\big{]}\Big{)},$ (130)
@@
  • the Chernoff bound in (128) is justified by $\widetilde{G}_{\mathcal{J}^{\prime},2}\sim\text{Bin}\big{(}(1-\rho)ne^{-\nu}(1+o(1)),\frac{\nu\ell(1+o(1))}{k}\big{)}$ (note also that $d<d_{0}$ implies $\rho d-(1-2\zeta)C<1-\rho$);
  • to obtain (130), we define $A=C(1-2\zeta)$ and relax the range of $d$ from $d\in\mathcal{D}$ to $d\geq\max\big{\{}0,\frac{A}{\rho}\big{\}}$, and we introduce the shorthand
    $g(C,\zeta,d,\rho)=\rho d\log d+\big{(}\rho d-A\big{)}\log\Big{(}\frac{\rho d-A}{1-\rho}\Big{)}+1-2\rho d+A.$ (131)
    Then, (130) follows from (129), by using $D(\frac{a\nu\ell}{k}\|\frac{\nu\ell}{k})=\frac{\nu\ell}{k}(a\log a-a+1+o(1))$, and noting that $g(C,\zeta,d,\rho)$ defined here coincides with (70) in the proof of converse bound for Bernoulli designs.
Thus, by the same reasoning as (71)–(73), we know that
$\min_{d\geq\max\{0,\frac{A}{\rho}\}}g(C,\zeta,d,\rho)=g(C,\zeta,d^{*},\rho):=f_{2}(C,\zeta,\rho),$ (132)
with $d^{*}$ being the same as in (73):
$d^{*}=\frac{A+\sqrt{A^{2}+4\rho(1-\rho)}}{2\rho}.$ (133)
Therefore, we obtain a bound on the term in (124) as
$\exp\Big{(}-\frac{\ell n\nu e^{-\nu}}{k}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}\Big{)},$ (134)
with $f_{2}(C,\zeta,\rho)$ being the same as in the proof of Bernoulli design converse bound; see (72).
Comparing the two bounds (134) and (125): We return to (129)–(130), in which we used $D(\frac{a\nu\ell}{k}\|\frac{\nu\ell}{k})=\frac{\nu\ell}{k}(a\log a-a+1+o(1))$ and established that
$\rho\cdot D\Big{(}\frac{d\nu\ell}{k}\big{\|}\frac{\nu\ell}{k}\Big{)}+(1-\rho)\cdot D\Big{(}\frac{[\rho d-(1-2\zeta)C]\nu\ell}{(1-\rho)k}\big{\|}\frac{\nu\ell}{k}\Big{)}=\frac{\nu\ell}{k}[g(C,\zeta,d,\rho)+o(1)].$ (135)
@@
$\frac{\ell n\nu e^{-\nu}}{k}f_{2}(C,\zeta,\rho)$ $=\frac{\ell n\nu e^{-\nu}}{k}\min_{d\geq\max\{0,\frac{A}{\rho}\}}g(C,\zeta,d,\rho)$ (136)
$\leq\frac{\ell n\nu e^{-\nu}}{k}g(C,\zeta,d_{0},\rho)$ (137)
$=ne^{-\nu}\Big{(}\rho D\Big{(}\frac{d_{0}\nu\ell}{k}\big{\|}\frac{\nu\ell}{k}\Big{)}+(1-\rho)D\Big{(}\frac{[\rho d_{0}-(1-2\zeta)C]\nu\ell}{(1-\rho)k}\big{\|}\frac{\nu\ell}{k}\Big{)}\Big{)}-o\Big{(}\frac{\ell n}{k}\Big{)}$ (138)
$=\rho ne^{-\nu}D\Big{(}\frac{d_{0}\nu\ell}{k}\big{\|}\frac{\nu\ell}{k}\Big{)}-o\Big{(}\frac{\ell n}{k}\Big{)}$ (139)
$=\rho ne^{-\nu}(1+o(1))D\Big{(}\frac{d_{0}\nu\ell}{k}\big{\|}\frac{\nu\ell}{k}\Big{)},$ (140)
where in (138) we substitute (135), and in (139) we substitute $d_{0}=\frac{(1-2\zeta)C+1-\rho}{\rho}$ and observe that the second term in (138) vanishes. Therefore, the bound (134) dominates the one in (125), and it follows that
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime}}\geq(1-2\zeta-o(1))\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq\exp\Big{(}-\frac{\ell n\nu e^{-\nu}}{k}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}\Big{)}.$ (141)
@@
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime}}\geq(1-2\zeta-o(1))\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq 1,$ (142)
@@
Combining Cases 1-2: Some algebra verifies that $f_{2}(C,\zeta,\rho)=0$ when $A=(1-2\zeta)C=2\rho-1$: Substituting this into (133) gives $d^{*}=1$, and further we have $f_{2}(C,\zeta,\rho)=g(C,\zeta,1,\rho)=0$ by using (131) and $A=2\rho-1$. Thus, we can define the continuous function
$\hat{f}_{2}(C,\zeta,\rho):=\begin{cases}f_{2}(C,\zeta,\rho)&\text{if }(1-2\zeta)C\geq 2\rho-1\\ 0&\text{if }(1-2\zeta)C<2\rho-1.\end{cases}$ (143)
@@
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime}}\geq(1-2\zeta-o(1))\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq\exp\Big{(}-\frac{\ell n\nu e^{-\nu}}{k}\big{(}\hat{f}_{2}(C,\zeta,\rho)+o(1)\big{)}\Big{)}$ (144)
@@
The condition for (C2): Now we can bound the probability of (C2) holding for some $(\ell,C,\zeta)$ as
@@
$\leq\binom{p}{\ell}\cdot\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime}}\geq(1-2\zeta-o(1))\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}+k^{-10\ell}$ (146)
$\leq\exp\Big{(}(1+o(1))\ell\log\frac{p}{\ell}-\frac{\ell n\nu e^{-\nu}}{k}\big{(}\hat{f}_{2}(C,\zeta,\rho)+o(1)\big{)}\Big{)}+k^{-10\ell},$ (147)
where (145) follows because under the condition (116) we have that (117) is a necessary condition for “(C2) holds for the given $(\ell,C,\zeta)$”, and in (146) we apply a union bound to account for the maximum over $\mathcal{J}^{\prime}$ in (117), also recalling that (116) holds with probability at least $1-k^{-10\ell}$.
Therefore, for given $(\ell,C,\zeta)$, we have identified the following condition for (C2) to hold (recall that our analysis on (C2) is on the event $\mathscr{A}_{1}$ (118), and in our conclusion below we make this explicit):
$\frac{\ell n\nu e^{-\nu}}{k}\big{[}\hat{f}_{2}(C,\zeta,\rho)+o(1)\big{]}\geq(1+\eta_{2})\ell\log\frac{p}{\ell},\leavevmode\nobreak\ \text{for some }\eta_{2}>0$ (148)
@@
$\quad\quad\leq\exp\Big{(}-\frac{\eta_{2}}{2}\ell\log\frac{p}{\ell}\Big{)}+k^{-10\ell}\leq\hat{P}_{\ell}:=\begin{cases}(\ell\log k)^{-5},\leavevmode\nobreak\ \text{if }1\leq\ell\leq\log k\\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ k^{-5},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }\log k<\ell\leq\frac{k}{\log k},\end{cases}$ (149)
since $\exp(-\frac{\eta_{2}}{2}\ell\log\frac{p}{\ell})+k^{-10\ell}$ is an upper bound on (147) when (148) holds. Recall also that $\mathbb{P}(\mathscr{A}_{1})=1-o(1)$.
Establishing the threshold
We are now ready to establish the threshold for $n$ above which restricted MLE has $o(1)$ probability of failing. For clarity, we pause to review our previous developments:
  • For given $\ell$, from (86)–(91), (C1) and (C2) hold for some feasible $(C,\zeta)\in(C_{0},\infty)\times[0,1]$ with probability at most $k^{-5}$.
  • For given $\ell$ and any feasible $(C,\zeta)$ with $C\leq C_{0}$, from (103)–(104), if
    $n\geq(1+\eta_{1})\frac{k\log\frac{k}{\ell}}{\nu e^{-\nu}[f_{1}(C,\zeta,\rho)+o(1)]}$ (150)
    holds for some $\eta_{1}>0$, then (C1) holds for the given $(\ell,C,\zeta)$ with probability at most $\hat{P}_{\ell}.$
  • For given $\ell$ and any feasible $(C,\zeta)$ with $C\leq C_{0}$, by (148)–(149), if
    $n\geq(1+\eta_{2})\frac{k\log\frac{p}{\ell}}{\nu e^{-\nu}[\hat{f}_{2}(C,\zeta,\rho)+o(1)]}$ (151)
@@
Bounding the failure probability for a fixed $\ell$: For a fixed $\ell$, we first consider the feasible $(C,\zeta)$ with $C\leq C_{0}$, and by $\frac{Cn\nu e^{-\nu}\ell}{k},\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}\in\mathbb{Z}$ there are no more than $O(\frac{\ell n}{k})\times O(\frac{\ell n}{k})=O((\ell\log k)^{2})$ possible choices of such $(C,\zeta)$. If it holds for some $\eta_{1}>0$ that
$n\geq(1+\eta_{1})\frac{k}{\nu e^{-\nu}}\max_{\begin{subarray}{c}0<C<C_{0}\\ 0<\zeta<1\end{subarray}}\min\Big{\{}\frac{\log\frac{k}{\ell}}{f_{1}(C,\zeta,\rho)},\frac{\log\frac{p}{\ell}}{\hat{f}_{2}(C,\zeta,\rho)}\Big{\}},$ (152)
@@
[…]
@@
Overall, for a fixed $\ell\in[1,\frac{k}{\log k}]$, if (152) holds for some $\eta_{1}>0$, then for the restricted MLE output $\widehat{S}^{\prime}$ (see (30)) we have
@@
$\leq\mathbb{P}\Big{(}\textbf{(C1)}\text{ and }\textbf{(C2)}\leavevmode\nobreak\ \text{hold for some feasible }(C,\zeta)\cap\,\mathscr{A}_{1}\,\text{ holds}\Big{)}$ (153)
$\leq k^{-5}+O\big{(}(\ell\log k)^{2}\big{)}\hat{P}_{\ell},$ (154)
@@
where (153) holds because “(C1) and (C2) hold for some feasible $(C,\zeta)$ (under the given $\ell$)” is a necessary condition for “$|s\setminus\widehat{S}^{\prime}|=\ell$” due to Lemma 3(b).

### 23 Oct 2024 v2

@@ Appendix D $-\ell$ Achievability Analysis for Bernoulli Designs
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime}}\geq(1-2\zeta-o(1))\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}\geq\frac{d_{0}\cdot\rho\nu\ell ne^{-\nu}}{k}\Big{)}$ (D.37)
$\quad\quad+\sum_{\begin{subarray}{c}0\leq d<d_{0}\\ \frac{d\rho\nu e^{-\nu}\ell n}{k}\in\mathbb{Z}\end{subarray}}\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}=\frac{d\cdot\rho\nu\ell ne^{-\nu}}{k}\Big{)}\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},2}\leq\big{[}\rho d-(1-2\zeta-o(1))C\big{]}\frac{n\nu e^{-\nu}\ell}{k}\Big{)}.$ (D.38)
To bound the terms in (D.37) and (D.38), we consider the following two cases.
@@
Bounding the term in (D.37): Consider conditioning on $(\mathbf{X}_{s},\mathbf{Y})$ and utilizing the randomness of $\mathbf{X}_{\mathcal{J}^{\prime}}$, and recall from (D.35) and (D.22) that we have $\widetilde{G}_{\mathcal{J}^{\prime},1}\sim\text{Bin}(\rho ne^{-\nu}(1+o(1)),\frac{\nu\ell(1+o(1))}{k})$. Since $d_{0}>1$, we can apply Chernoff bound ((A.1) in Lemma A.1) to bound the term in (D.37) as
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}\geq\frac{d_{0}\cdot\rho\nu\ell ne^{-\nu}}{k}\Big{)}$ $\leq\exp\left(-\rho ne^{-\nu}(1+o(1))D\Big{(}\frac{d_{0}\ell\nu}{k}\big{\|}\frac{\ell\nu}{k}\Big{)}\right).$ (D.39)
Bounding the term in (D.38): First observe that $\mathbb{P}(\widetilde{G}_{\mathcal{J}^{\prime},2}\leq[\rho d-(1-2\zeta-o(1))C]\frac{n\nu e^{-\nu}\ell}{k})=0$ if $d<\frac{C(1-2\zeta)}{\rho}-o(1)$, and so we can further restrict the summation over $d$ in (D.38) to the range
$\mathcal{D}:=\Big{\{}\max\Big{\{}0,\frac{C(1-2\zeta)}{\rho}\Big{\}}\leq d<d_{0}:\frac{d\rho\nu e^{-\nu}\ell n}{k}\in\mathbb{Z}\Big{\}}$ (D.40)
up to a $o(1)$ term that has no impact on our subsequent analysis. Since $d_{0}=O(1)$, the term in (D.38) involves no more than $O(\frac{\ell n}{k})$ summands, so using the Chernoff bound (see (A.1) in Lemma A.1) we can bound it as
$O\Big{(}\frac{\ell n}{k}\Big{)}\max_{d\in\mathcal{D}}\Big{\{}\exp\Big{(}-\rho ne^{-\nu}(1+o(1))\cdot D\Big{(}\frac{d\nu\ell(1+o(1))}{k}\big{\|}\frac{\nu\ell(1+o(1))}{k}\Big{)}\Big{)}$ (D.41)
$\quad\quad\quad\cdot\exp\Big{(}-(1-\rho)ne^{-\nu}(1+o(1))\cdot D\Big{(}\frac{[\rho d-(1-2\zeta)C]\nu\ell(1+o(1))}{(1-\rho)k}\big{\|}\frac{\nu\ell(1+o(1))}{k}\Big{)}\Big{)}\Big{\}}$ (D.42)
$=\exp\Big{(}-ne^{-\nu}(1+o(1))\min_{d\in\mathcal{D}}\Big{[}\rho\cdot D\Big{(}\frac{d\nu\ell}{k}\big{\|}\frac{\nu\ell}{k}\Big{)}+(1-\rho)D\Big{(}\frac{[\rho d-(1-2\zeta)C]\nu\ell}{(1-\rho)k}\big{\|}\frac{\nu\ell}{k}\Big{)}\Big{]}\Big{)}$ (D.43)
$\leq\exp\Big{(}-\frac{\ell n\nu e^{-\nu}}{k}\big{[}\min_{d\geq\max\{0,\frac{A}{\rho}\}}g(C,\zeta,d,\rho)+o(1)\big{]}\Big{)},$ (D.44)
@@
  • the Chernoff bound in (D.42) is justified by $\widetilde{G}_{\mathcal{J}^{\prime},2}\sim\text{Bin}\big{(}(1-\rho)ne^{-\nu}(1+o(1)),\frac{\nu\ell(1+o(1))}{k}\big{)}$ (note also that $d<d_{0}$ implies $\rho d-(1-2\zeta)C<1-\rho$);
  • to obtain (D.44), we define $A=C(1-2\zeta)$ and relax the range of $d$ from $d\in\mathcal{D}$ to $d\geq\max\big{\{}0,\frac{A}{\rho}\big{\}}$, and we introduce the shorthand
    $g(C,\zeta,d,\rho)=\rho d\log d+\big{(}\rho d-A\big{)}\log\Big{(}\frac{\rho d-A}{1-\rho}\Big{)}+1-2\rho d+A.$ (D.45)
    Then, (D.44) follows from (D.43), by using $D(\frac{a\nu\ell}{k}\|\frac{\nu\ell}{k})=\frac{\nu\ell}{k}(a\log a-a+1+o(1))$, and noting that $g(C,\zeta,d,\rho)$ defined here coincides with (C.34) in the proof of converse bound for Bernoulli designs.
Thus, by the same reasoning as (C.35)–(C.37), we know that
$\min_{d\geq\max\{0,\frac{A}{\rho}\}}g(C,\zeta,d,\rho)=g(C,\zeta,d^{*},\rho):=f_{2}(C,\zeta,\rho),$ (D.46)
with $d^{*}$ being the same as in (C.37):
$d^{*}=\frac{A+\sqrt{A^{2}+4\rho(1-\rho)}}{2\rho}.$ (D.47)
Therefore, we obtain a bound on the term in (D.38) as
$\exp\Big{(}-\frac{\ell n\nu e^{-\nu}}{k}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}\Big{)},$ (D.48)
with $f_{2}(C,\zeta,\rho)$ being the same as in the proof of Bernoulli design converse bound; see (C.36).
Comparing the two bounds (D.48) and (D.39): We return to (D.43)–(D.44), in which we used $D(\frac{a\nu\ell}{k}\|\frac{\nu\ell}{k})=\frac{\nu\ell}{k}(a\log a-a+1+o(1))$ and established that
$\rho\cdot D\Big{(}\frac{d\nu\ell}{k}\big{\|}\frac{\nu\ell}{k}\Big{)}+(1-\rho)\cdot D\Big{(}\frac{[\rho d-(1-2\zeta)C]\nu\ell}{(1-\rho)k}\big{\|}\frac{\nu\ell}{k}\Big{)}=\frac{\nu\ell}{k}[g(C,\zeta,d,\rho)+o(1)].$ (D.49)
@@
$\frac{\ell n\nu e^{-\nu}}{k}f_{2}(C,\zeta,\rho)$ $=\frac{\ell n\nu e^{-\nu}}{k}\min_{d\geq\max\{0,\frac{A}{\rho}\}}g(C,\zeta,d,\rho)$ (D.50)
$\leq\frac{\ell n\nu e^{-\nu}}{k}g(C,\zeta,d_{0},\rho)$ (D.51)
$=ne^{-\nu}\Big{(}\rho D\Big{(}\frac{d_{0}\nu\ell}{k}\big{\|}\frac{\nu\ell}{k}\Big{)}+(1-\rho)D\Big{(}\frac{[\rho d_{0}-(1-2\zeta)C]\nu\ell}{(1-\rho)k}\big{\|}\frac{\nu\ell}{k}\Big{)}\Big{)}-o\Big{(}\frac{\ell n}{k}\Big{)}$ (D.52)
$=\rho ne^{-\nu}D\Big{(}\frac{d_{0}\nu\ell}{k}\big{\|}\frac{\nu\ell}{k}\Big{)}-o\Big{(}\frac{\ell n}{k}\Big{)}$ (D.53)
$=\rho ne^{-\nu}(1+o(1))D\Big{(}\frac{d_{0}\nu\ell}{k}\big{\|}\frac{\nu\ell}{k}\Big{)},$ (D.54)
where in (D.52) we substitute (D.49), and in (D.53) we substitute $d_{0}=\frac{(1-2\zeta)C+1-\rho}{\rho}$ and observe that the second term in (D.52) vanishes. Therefore, the bound (D.48) dominates the one in (D.39), and it follows that
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime}}\geq(1-2\zeta-o(1))\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq\exp\Big{(}-\frac{\ell n\nu e^{-\nu}}{k}\big{(}f_{2}(C,\zeta,\rho)+o(1)\big{)}\Big{)}.$ (D.55)
@@
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime}}\geq(1-2\zeta-o(1))\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq 1,$ (D.56)
@@
Combining Cases 1-2: Some algebra verifies that $f_{2}(C,\zeta,\rho)=0$ when $A=(1-2\zeta)C=2\rho-1$: Substituting this into (D.47) gives $d^{*}=1$, and further we have $f_{2}(C,\zeta,\rho)=g(C,\zeta,1,\rho)=0$ by using (D.45) and $A=2\rho-1$. Thus, we can define the continuous function
$\hat{f}_{2}(C,\zeta,\rho):=\begin{cases}f_{2}(C,\zeta,\rho)&\text{if }(1-2\zeta)C\geq 2\rho-1\\ 0&\text{if }(1-2\zeta)C<2\rho-1.\end{cases}$ (D.57)
@@
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime}}\geq(1-2\zeta-o(1))\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq\exp\Big{(}-\frac{\ell n\nu e^{-\nu}}{k}\big{(}\hat{f}_{2}(C,\zeta,\rho)+o(1)\big{)}\Big{)}$ (D.58)
@@
The condition for (C2): We now bound the probability of (C2) holding for some $(\ell,C,\zeta)$ as
@@
$\leq\binom{p}{\ell}\cdot\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime}}\geq(1-2\zeta-o(1))\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}+k^{-10\ell}$ (D.60)
$\leq\exp\Big{(}(1+o(1))\ell\log\frac{p}{\ell}-\frac{\ell n\nu e^{-\nu}}{k}\big{(}\hat{f}_{2}(C,\zeta,\rho)+o(1)\big{)}\Big{)}+k^{-10\ell},$ (D.61)
where (D.59) follows because under the condition (D.32) we have that (D.33) is a necessary condition for “(C2) holds for the given $(\ell,C,\zeta)$”, and in (D.60) we apply a union bound to account for the maximum over $\mathcal{J}^{\prime}$ in (D.33), also recalling that (D.32) holds with probability at least $1-k^{-10\ell}$.
Therefore, for given $(\ell,C,\zeta)$, we have shown the following implication for any $\eta_{2}>0$ and sufficiently large $p$ (recall that our analysis of (C2) is on the event $\mathscr{A}_{1}$ (D.34), and in our conclusion below we make this explicit):
$\frac{\ell n\nu e^{-\nu}}{k}\big{[}\hat{f}_{2}(C,\zeta,\rho)+o(1)\big{]}\geq(1+\eta_{2})\ell\log\frac{p}{\ell}$ (D.62)
($\Longrightarrow\leavevmode\nobreak\ \mathbb{P}\Big{(}\big{\{}\textbf{(C2)}\text{ holds for }(\ell,C,\zeta)\big{\}}\cap\mathscr{A}_{1}\Big{)}$ $\leq\exp\Big{(}-\frac{\eta_{2}}{2}\ell\log\frac{p}{\ell}\Big{)}+k^{-10\ell}$ (D.63)
$\leq\hat{P}_{\ell}:=\begin{cases}(\ell\log k)^{-5},\leavevmode\nobreak\ \text{if }1\leq\ell\leq\log k\\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ k^{-5},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }\log k<\ell\leq\frac{k}{\log k},\end{cases}$ (D.64)
since $\exp(-\frac{\eta_{2}}{2}\ell\log\frac{p}{\ell})+k^{-10\ell}$ is an upper bound on (D.61) when (D.62) holds. Recall also that $\mathbb{P}(\mathscr{A}_{1})=1-o(1)$.
Establishing the threshold
We are now ready to establish the threshold for $n$ above which restricted MLE has $o(1)$ probability of failing. We first pause to review our previous developments:
  • For given $\ell$, from (D.3)–(D.8), (C1) and (C2) hold for some feasible $(C,\zeta)\in(C_{0},\infty)\times[0,1]$ with probability at most $k^{-5}$.
  • For given $\ell$ and any feasible $(C,\zeta)$ with $C\leq C_{0}$, from (D.20)–(D.21), if
    $n\geq(1+\eta_{1})\frac{k\log\frac{k}{\ell}}{\nu e^{-\nu}[f_{1}(C,\zeta,\rho)+o(1)]}$ (D.65)
@@
  • For given $\ell$ and any feasible $(C,\zeta)$ with $C\leq C_{0}$, by (D.62)–(D.64), if
    $n\geq(1+\eta_{2})\frac{k\log\frac{p}{\ell}}{\nu e^{-\nu}[\hat{f}_{2}(C,\zeta,\rho)+o(1)]}$ (D.66)
@@
Bounding the failure probability for a fixed $\ell$: For a fixed $\ell$, we first consider the feasible $(C,\zeta)$ with $C\leq C_{0}$, and by $\frac{Cn\nu e^{-\nu}\ell}{k},\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}\in\mathbb{Z}$ there are no more than $O(\frac{\ell n}{k})\times O(\frac{\ell n}{k})=O((\ell\log k)^{2})$ possible choices of such $(C,\zeta)$. (Recall from Section A.2 that $n=\Theta(k\log k)$.) If it holds for some $\eta_{1}>0$ that
$n\geq(1+\eta_{1})\frac{k}{\nu e^{-\nu}}\max_{\begin{subarray}{c}0<C<C_{0}\\ 0<\zeta<1\end{subarray}}\min\Big{\{}\frac{\log\frac{k}{\ell}}{f_{1}(C,\zeta,\rho)},\frac{\log\frac{p}{\ell}}{\hat{f}_{2}(C,\zeta,\rho)}\Big{\}},$ (D.67)
@@
((C1) and (C2) hold for $\mathbb{P}\big{(}(\textbf{C1})\text{ and }(\textbf{C2})\text{ hold for }(C,\zeta)\,\cap\,\mathscr{A}_{1}\text{ holds}\big{)}\leq\hat{P}_{\ell}.$
@@
Overall, for a fixed $\ell\in[1,\frac{k}{\log k}]$, if (D.67) holds for some $\eta_{1}>0$, then for the restricted MLE output $\widehat{S}^{\prime}$ (see (B.2)) we have
@@
$\leq\mathbb{P}\Big{(}\textbf{(C1)}\text{ and }\textbf{(C2)}\leavevmode\nobreak\ \text{hold for some feasible }(C,\zeta)\cap\,\mathscr{A}_{1}\,\text{ holds}\Big{)}$ (D.68)
$\leq k^{-5}+O\big{(}(\ell\log k)^{2}\big{)}\hat{P}_{\ell},$ (D.69)
@@
where (D.68) holds because “(C1) and (C2) hold for some feasible $(C,\zeta)$ (under the given $\ell$)” is a necessary condition for “$|s\setminus\widehat{S}^{\prime}|=\ell$” due to Lemma B.1(b).